RENNER-TELLER COUPLING IN H_{2}S+: PARTITIONING THE ROVIBRONIC AND SPINORBIT COU-PLING HAMILTONIAN

\underline{G. DUXBURY}, Department of Physics, SUPA, John Anderson Building, University of Strathclyde, 107 Rottenrow, Glasgow G4 0NG, Scotland, UK; Ch. JUNGEN, LAC, 1 Laboratoire Aimé Cotton du CNRS, Université de Paris-Sud, 91405 Orsay, France; A. ALIJAH, GSMA, UMR CNRS 6089, Université de Reims Champagne-Ardenne,B.P. 1039, 51687 Reims Cedex 2, France.

The stretch-bender reference-frame^{a} was developed to allow the separation of large amplitude bending motion and symmetric stretching. It has been used to calculate vibrational resonances, the effects of spin-orbit coupling, and of overall rotation. Its use allows the block factorisation of the Renner-Teller interaction matrix. We wish to show the utility of this approach when two different approaches by Dixon and Duxbury,^{b} and Jungen and Merer^{c}, are used to minimise the effects of the large amplitude bending upon the Renner-Teller interaction. It also allows the effects of large amplitude motion on the rotational structure to be calculated, including the switchover from bent to linear behaviour.

aJ. Chem. Phys. 108, 2336 (1998) and J. Mol. Spectrosc. 211, 7 (2002)

bMolec. Phys. 43, 255 (1981)

cMolec. Phys. 40, 1 (1981)